# Intrinsic Dimensionality as a Model-Free Measure of Class Imbalance

### v1

@@ Abstract
Imbalance in classification tasks is commonly quantified by the cardinalities of examples across classes. This, however, disregards the presence of redundant examples and inherent differences in the learning difficulties of classes. Alternatively, one can use complex measures such as training loss and uncertainty, which, however, depend on training a machine learning model. Our paper proposes using data Intrinsic Dimensionality (ID) as an easy-to-compute, model-free measure of imbalance that can be seamlessly incorporated into various imbalance mitigation methods. Our results across five different datasets with a diverse range of imbalance ratios show that ID consistently outperforms cardinality-based re-weighting and re-sampling techniques used in the literature. Moreover, we show that combining ID with cardinality can further improve performance. Code: https://github.com/cagries/IDIM.
keywords:
intrinsic dimension , long-tailed visual recognition , class imbalance , long-tailed learning
journal: Preprint submitted to arXiv
[…]
[ceng]organization=Department of Computer Engineering, addressline=Middle East Technical University, city=Ankara, postcode=06810, state=Ankara, country=Turkey
[…]
@@ 1 Introduction
Figure 2: Estimated IDs for CIFAR10 classes for different sample counts. ID can capture inherent differences among classes and is relatively robust against sample count, as exemplified here with the CIFAR10 dataset, subsampled at different sample counts per class.
@@ 2.1 Long-tailed Visual Recognition
Long-tailed recognition approaches mitigate the effects of imbalance relying on sample counts or hardness under the following main categories [2, 1, 20]: Resampling methods [7, 6], loss-reweighting methods [8, 9, 21, 22], margin (logit) adjustment methods [23, 24] and other methods [25, 26, 27]. For a detailed review, see A.1 and, for formal definitions of the common approaches, see Section 5.
@@ 2.2 Intrinsic Dimensionality
Intrinsic dimensionality analysis focuses on quantifying the intrinsic qualities of data. As a general definition, the intrinsic dimension is the minimum number of parameters required to represent a data distribution with minimal loss of generality [11]. ID is closely related to the manifold hypothesis [14], which posits that high dimensional natural image datasets lie on a lower dimensional manifold, enabling learning. In this context, ID refers to the number of dimensions of this lower-dimensional manifold. The ID of a dataset can be estimated globally (i.e., for the entire dataset), or locally (around a local neighborhood), depending on the scope and the estimation method. ID estimation methods in the literature can be studied in two general categories: Projective methods [28, 29, 30, 31] and geometric (topological) methods [32, 33, 34]. For the sake of space, we provide a further review of these methods in A.2.
@@ 4 ID as a Measure of Class Imbalance
ID estimation is robust against correlation. We also show that ID estimation is relatively robust against dimension correlation. See D.4 for this analysis.
@@ 6 Experiments and Results
Datasets. We evaluate our method through the CIFAR-LT(10/100), Places-LT and ImageNet-LT datasets that are commonly used in the long-tailed visual recognition literature, as well as the semantically imbalanced SVCI-20 dataset. For details about the datasets, see C.1.
@@
Training & Implementation Details are provided in C.2.
@@
Table 1: Exp. 1: Top-1 accuracy results for plug-and-play reweighting and resampling methods for ResNet-32 on long-tailed CIFAR datasets. [2S]: Two-stage methods.
@@ 6.1 Experiment 1: Model-Free Approaches
Re-sampling and Re-weighting. The results on the CIFAR-LT datasets in Table 1 show that both ID-based resampling and ID-based reweighting methods outperform alternative methods. Furthermore, we observe that our method keeps its effectiveness under increasing data imbalance, in our detailed analysis of one-stage methods in D.
@@
Margin Adjustment. Aside from reweighting and resampling, we also experiment with integrating ID into margin-based methods on a case-by-case basis in Table 3 (additional results on CIFAR-100 can be found in D). In this case, ID-based mitigation performs better in some cases while providing second-best results in others.
@@
BoT + ID [Ours] 43.39 42.92
Table 2: Exp. 1: Top-1 accuracy results for plug-and-play reweighting and resampling methods on Places-LT and ImageNet-LT. Two-stage methods marked with [2S].
@@
Table 3: Exp. 1: Top-1 accuracy results for simple plug-and-play margin-based methods on long-tailed CIFAR-10. N.: normalized. L: Learned epsilon. Log. Adj.: Logit Adjustment. See D for the results on CIFAR-100-LT.
@@ 6.2 Experiment 2: SOTA Comparison
Table 4: Exp. 2: Top-1 accuracy results for state-of-the-art methods on long-tailed CIFAR datasets with a ResNet-32 backbone. denotes results reproduced using the code released by the authors.
@@
PB-ID [Ours] - 42.6
BoT + ID [Ours] - 43.4
Table 5: Exp. 2: Top-1 accuracy results for state-of-the-art methods on Places-LT with a ResNet-152 backbone.
@@ 6.3 Experiment 3: ID and Semantic Imbalance
As shown in Figure 5, the individual ID estimates of CIFAR-10 classes are much higher than SVHN classes, indicating a clear separation of CIFAR-10 classes from SVHN classes within the SVCI-20 dataset when using ID-based model training. This shows that our ID-based approach does capture imbalance among classes where cardinality-based methods fail. Individual ID values for SVCI-20 classes can additionally be found in D.4.
@@
Table 6: Exp. 3: Top-1 balanced accuracy values for plug-and-play resampling and reweighting methods on the balanced SVCI-20 dataset. NN: Number of samples per class. [*]: [4, 8, 7, 6].
Figure 5: Exp. 3: Histogram of estimated ID values of the two semantically-different types of classes in the SVCI-20 dataset. See D.4 for all ID values.
6.4 Experiment 4: Ablation Analysis
We now evaluate our ID-based approach with different ID estimators: FisherS [29], maximum likelihood estimator (MLE) [32] with asymptotic correction [53], and Tight Locality Estimation (TLE) [33]. For comparison on CIFAR-LT, we use the GLMC [43] and SURE + re-weighting methods [51] (Table 8), and for ImageNet-LT, we use BCL [44] – see Appendix for all results.
Summary. Overall, we see that our approach works with other ID estimators and that there is no single ID estimator which is best for all methods. With SURE + ID and BCL + ID, our method equipped with the FisherS estimator works best, while with GLMC + ID we observe that our method with TLE and MLE provides better results.
@@ 6.4 Experiment 4: Ablation Analysis
Table 7: Exp. 2: Top-1 accuracy results for state-of-the-art methods on ImageNet-LT under 100 and 180 epochs. denotes methods that use RandAugment [58] augmentation. 3E: 3 experts. EP: Epochs.
@@
Table 8: Exp. 4: Top-1 accuracy results for ID-based methods with different ID estimators on CIFAR-LT. denotes results reproduced using the code released by the authors.
7 Conclusion
@@ Acknowledgments
Appendix A Extended Literature
A.1 Literature on Long-Tailed Visual Recognition

### v2

@@ Abstract
Imbalance in classification tasks is commonly quantified by the cardinalities of examples across classes. This, however, disregards the presence of redundant examples and inherent differences in the learning difficulties of classes. Alternatively, one can use complex measures such as training loss and uncertainty, which, however, depend on training a machine learning model. Our paper proposes using data Intrinsic Dimensionality (ID) as an easy-to-compute, model-free measure of imbalance that can be seamlessly incorporated into various imbalance mitigation methods. Our results across five different datasets with a diverse range of imbalance ratios show that ID consistently outperforms cardinality-based re-weighting and re-sampling techniques used in the literature. Moreover, we show that combining ID with cardinality can further improve performance. Our code and models are available at https://github.com/cagries/IDIM.
footnotetext:
To appear in Neurocomputing.
Keywords intrinsic dimensionality $\cdot$ class imbalance $\cdot$ long-tailed learning
@@ 1 Introduction
Figure 2: Estimated IDs for CIFAR-10 classes for different sample counts. ID can capture inherent differences among classes and is relatively robust against sample count, as exemplified here with the CIFAR-10 dataset, subsampled at different sample counts per class.
@@ 2.1 Long-tailed Visual Recognition
Long-tailed recognition approaches mitigate the effects of imbalance relying on sample counts or hardness under the following main categories [71, 75, 76]: Resampling methods [38, 50], loss-reweighting methods [37, 13, 27, 73], margin (logit) adjustment methods [57, 34] and other methods [18, 3, 74]. For a detailed review, see Appendix B.1 and, for formal definitions of the common approaches, see Section 5.
@@ 2.2 Intrinsic Dimensionality
Intrinsic dimensionality analysis focuses on quantifying the intrinsic qualities of data. As a general definition, the intrinsic dimension is the minimum number of parameters required to represent a data distribution with minimal loss of generality [30]. ID is closely related to the manifold hypothesis [31], which posits that high dimensional natural image datasets lie on a lower dimensional manifold, enabling learning. In this context, ID refers to the number of dimensions of this lower-dimensional manifold. The ID of a dataset can be estimated globally (i.e., for the entire dataset), or locally (around a local neighborhood), depending on the scope and the estimation method. ID estimation methods in the literature can be studied in two general categories: Projective methods [8, 2, 59, 65] and geometric (topological) methods [39, 4, 25]. For the sake of space, we provide a further review of these methods in Appendix B.2.
@@ 4 ID as a Measure of Class Imbalance
ID estimation is robust against correlation. We also show that ID estimation is relatively robust against dimension correlation. See Appendix E.3 for this analysis.
4.1 Theoretical Motivation
To theoretically motivate our use of class-wise ID as a suitable measure of imbalance, we follow the class-wise manifold analysis work by Narayanan & Niyogi [53]. Narayanan & Niyogi proved that the number of balls of radius $\epsilon$ required to cover the space of smooth boundaries on a $d$-dimensional manifold is related to the manifold’s volume and intrinsic dimension and to sample complexity.
Let the total support of the data $X$ be the union of $K$ disjoint compact Riemannian sub-manifolds, where each manifold $\mathcal{M}_{k}$ corresponds to a class indexed by $k$:
$X=\bigcup_{k=1}^{K}\mathcal{M}_{k}.$ (6)
Each class manifold $\mathcal{M}_{k}$ has its own distinct intrinsic dimension $d_{k}$. Moreover, let $R_{k}(f)=\mathbb{E}_{x\sim\mathcal{M}_{k}}[\mathbb{I}(f(x)\neq k)]$ denote the generalization error specifically on class $k$.
Theorem 1 (Class-wise Intrinsic Dimension and Sample Complexity).
The number of samples $n_{k}$ for class $k$ required to bound the generalization error $R_{k}(f)\leq\epsilon$ with high probability is:
$n_{k}\gtrsim\Omega\left(\left(\frac{1}{\epsilon}\right)^{d_{k}}\right).$ (7)
Proof.
See Appendix A for a sketch of the proof. ∎
This theorem suggests that a class with low ID ($d_{k}$ small) is easy to learn even with few samples. On the other hand, a class with high ID ($d_{k}$ large) requires exponentially more samples. If we calculated a single “global” ID for the whole dataset, we would average out these differences, potentially masking the difficulty of a complex minority class.
Motivated by the theoretical link between class-wise ID and sample complexity, we propose using class-wise ID ($d_{k}$) as a measure of class imbalance, which we can use in place of class cardinality in designing sampling-based, loss-reweighting-based, or logit-adjustment-based imbalance mitigation strategies. However, if a class is mostly noise, lacking sufficient structure, its ID estimation will be high, which can negatively affect the imbalance mitigation strategy. This, on the other hand, is assumed to be rare in practice.
@@ 6 Experiments and Results
Datasets. We evaluate our method through the CIFAR-10-LT and CIFAR-100-LT, Places-LT and ImageNet-LT datasets that are commonly used in the long-tailed visual recognition literature, as well as the semantically imbalanced SVCI-20 dataset. For details about the datasets, see Appendix D.1.
@@
Training & Implementation Details are provided in Appendix D.2.
@@
Table 1: Experiment 1: Top-1 accuracy results for plug-and-play reweighting and resampling methods for ResNet-32 on long-tailed CIFAR datasets. [2S]: Two-stage methods.
@@ 6.1 Experiment 1: Model-Free Approaches
Re-sampling and Re-weighting. The results on the CIFAR-LT datasets in Table 1 show that both ID-based resampling and ID-based reweighting methods outperform alternative methods. Furthermore, we observe that our method keeps its effectiveness under increasing data imbalance, in our detailed analysis of one-stage methods in Appendix E.
@@
Margin Adjustment. Aside from reweighting and resampling, we also experiment with integrating ID into margin-based methods on a case-by-case basis in Table 3 (additional results on CIFAR-100 can be found in Appendix E). In this case, ID-based mitigation performs better in some cases while providing second-best results in others.
@@
BoT + ID [Ours] 43.39 $\pm$ 0.09 42.92 $\pm$0.11
Table 2: Experiment 1: Top-1 accuracy results for plug-and-play reweighting and resampling methods on Places-LT and ImageNet-LT. Two-stage methods marked with [2S].
@@
Table 3: Experiment 1: Top-1 accuracy results for simple plug-and-play margin-based methods on long-tailed CIFAR-10. N.: normalized. L: Learned epsilon. Log. Adj.: Logit Adjustment. See Appendix E for the results on CIFAR-100-LT.
@@ 6.2 Experiment 2: SOTA Comparison
Table 4: Experiment 2: Top-1 accuracy results for state-of-the-art methods on long-tailed CIFAR datasets with a ResNet-32 backbone. denotes results reproduced using the code released by the authors.
@@
PB-ID [Ours] - 42.6 $\pm$ 0.15
BoT + ID [Ours] - 43.4 $\pm$ 0.09
Table 5: Experiment 2: Top-1 accuracy results for state-of-the-art methods on Places-LT with a ResNet-152 backbone.
@@ 6.3 Experiment 3: ID and Semantic Imbalance
As shown in Fig. 5, the individual ID estimates of CIFAR-10 classes are much higher than SVHN classes, indicating a clear separation of CIFAR-10 classes from SVHN classes within the SVCI-20 dataset when using ID-based model training. This shows that our ID-based approach does capture imbalance among classes where cardinality-based methods fail. Individual ID values for SVCI-20 classes can additionally be found in Appendix E.3.
@@
Table 6: Experiment 3: Top-1 balanced accuracy values for plug-and-play resampling and reweighting methods on the balanced SVCI-20 dataset. NN: Number of samples per class.
Figure 5: Experiment 3: Histogram of estimated ID values of the two semantically-different types of classes in the SVCI-20 dataset. See Appendix E.3 for all ID values.
6.4 Experiment 4: Ablation on the Choice of ID Estimator
Considering that there are many ways to estimate the intrinsic dimensionality of a class, we now evaluate our ID-based approach with different ID estimators: FisherS [2], maximum likelihood estimator (MLE) [39] with asymptotic correction [49], and Tight Locality Estimation (TLE) [4]. For comparison on CIFAR-LT, we use the GLMC [24] and SURE + re-weighting methods [42], and for ImageNet-LT, we use BCL [81]. Table 8 shows the results for our method paired up with the GLMC and SURE on CIFAR-LT datasets, while Table 9 indicates that for ImageNet-LT, the FisherS estimator performs better than alternative estimators when paired up with the BCL [81] method.
In general, our results indicate that there is no singular optimal choice of ID estimator for all cases, and that the optimal estimator may be different for a particular dataset and a particular mitigation method. However, based on its all-around performance improvements on multiple datasets and its demonstrated empirical stability for natural image datasets, we suggest that our ID-based method paired with the FisherS estimator provides the most stable and consistent results.
Summary. Overall, we see that our approach works with other ID estimators and that there is no single ID estimator which is best for all methods. With SURE + ID and BCL + ID, our method equipped with the FisherS estimator works best, while with GLMC + ID we observe that TLE and MLE provides better results. Considering all options, our method with the FisherS estimator provides the most consistent improvements.
@@ 6.4 Experiment 4: Ablation on the Choice of ID Estimator
Table 7: Experiment 2: Top-1 accuracy results for state-of-the-art methods on ImageNet-LT under 100 and 180 epochs. denotes methods that use RandAugment [17] augmentation. 3E: 3 experts. EP: Epochs.
@@
Table 8: Experiment 4: Top-1 accuracy results for ID-based methods with different ID estimators on CIFAR-LT. denotes results reproduced using the code released by the authors.
Dataset: ImageNet-LT
BCL [81] 56.1
BCL + ID (w/ MLE) 54.4
BCL + ID (w/ TLE) 55.5
BCL + ID (w/ FisherS) 56.5
6.5 Experiment 5: Failure Cases with ID-based Mitigation
We also briefly explore some cases where our ID-based method may fail to be effective in imbalance mitigation. One possible failure case for our methods arises when the estimated ID fails to be informative for a particular class. One way to analyze this is to reverse or shuffle the estimated ID values for each class in the dataset. Table 10 shows the result of mitigation with uninformative ID values on CIFAR-10, where the results trail behind standard CE-based training.
Dataset: CIFAR-10-LT CIFAR-100-LT
CE 71.60 38.90
CSCE [37] 74.69 41.08
ID 76.44 41.65
ID (Reversed) 71.55 39.33
ID (Shuffled) 71.52 $\pm$ 0.85 38.84 $\pm$ 0.61
Table 10: Experiment 5: Top-1 accuracy results with uninformative ID configurations for CIFAR-LT datasets with an imbalance ratio of 100.
Figure 6: Experiment 5: ID estimates on CIFAR-10 with low sample sizes.
6.5.1 Fallback Strategies for ID-based Mitigation
In this section, we briefly consider two fallback strategies for ID-based imbalance mitigation in the case of noisy ID estimates.
First, as a simpler construction to handle noisy ID estimates, we consider a combination strategy for tail classes in the dataset in the following manner: We modify the class-wise ID estimates $d_{c}$ as $\tilde{d_{c}}$, which is defined as:
$\tilde{d_{c}}=\lambda_{c}d_{c}+(1-\lambda_{c})\bar{d},$ (18)
where $d_{c}$ is the estimated ID for a class, $\bar{d}=\frac{1}{\mathcal{|C|}}\sum_{c}d_{c}$ represents the mean ID of all classes in the dataset, and $\lambda_{c}$ is a scaling parameter that scales the weighting depending on the cardinality of the class $c$. We consider two schedules for $\lambda_{c}$, a linear scaling which we define as
$\lambda_{c}^{\mathrm{linear}}=\frac{N_{c}-N_{\min}}{N_{\max}-N_{\min}},$ (19)
and a log-scaled version defined as
$\lambda_{c}^{\mathrm{log}}=\frac{\log N_{c}-\log N_{\min}}{\log N_{\max}-\log N_{\min}}.$ (20)
This mean-based formulation of $\tilde{d_{c}}$ reduces the variance of ID estimates for tail classes by moving $\tilde{d_{c}}$ towards the mean ID. Fig. 7 shows an example of combined ID values for CIFAR-100-LT.
Figure 7: Experiment 5: ID estimates with mean-based fallback strategies for CIFAR-100-LT with an imbalance ratio of $\rho=50$.
Second, we adopt a gating-based approach based on the following strategy: for each class, we construct a bootstrap distribution from repeated ID estimates with random 80% subsampling of their original samples. We then calculate confidence intervals for the ID estimates of each class, and gate our mitigation strategy with the following rule: if the CI range of the estimated ID is over a fixed value, we replace the ID-based $d_{c}$ weighting for that class with the cardinality-based weighting [37] instead. Fig. 8 shows an example of bootstrapped ID for CIFAR-10-LT with an imbalance ratio of $\rho=100$.
Figure 8: Experiment 5: ID class-wise estimates ($d_{c}$) along with their bootstrapped ID ranges, for CIFAR-10-LT with an imbalance ratio of $\rho=50$.
Table 11 shows the results of training model on CIFAR-LT datasets with these two fallback strategies. In particular, while ID with cardinality-based gating does not seem to have much effect, ID with mean fallback seems to slightly improve performance for the CIFAR-100-LT dataset with an imbalance ratio of 100. We hypothesize that gating might not be effective due to the mixing of ID-based and cardinality-based weights, while mean fallback may be useful for datasets with many tails of low sample sizes.
Dataset: CIFAR-10-LT CIFAR-100-LT
Imbalance Ratio ($\rho$): 100 50 100 50
ID 76.4 80.8 41.7 46.2
ID w/ mean fallback ($\lambda_{c}^{\mathrm{linear}}$) 75.9 $\pm$ 0.7 80.9 $\pm$ 0.3 42.1 $\pm$ 0.2 45.9 $\pm$ 0.3
ID w/ mean fallback ($\lambda_{c}^{\mathrm{log}}$) 75.8 $\pm$ 0.5 80.8 $\pm$ 0.4 41.5 $\pm$ 0.4 45.8 $\pm$ 0.5
ID w/ gating fallback 75.3 $\pm$ 0.5 80.5 $\pm$ 0.4 40.2 $\pm$ 0.3 45.8 $\pm$ 0.6
Table 11: Experiment 5: Top-1 accuracy results for fallback strategies for ID-based re-weighting.
Summary. Similar to other mitigation methods in the literature, ID-based mitigation also includes possible failure cases, particularly in the case of classes with extremely low sample sizes which can result in noisy ID estimates. Using uninformative or high-variance ID estimates can shift the model towards focusing on the wrong classes, and lead to suboptimal training and performance.
7 Conclusion and Discussion
7.1 Conclusion
@@
7.2 Practical Guidance
Considering all experiments about ID-based mitigation throughout this section, we suggest the following setup for ID-based mitigation.
Estimating ID. Considering all available choices for ID estimator, we suggest using the FisherS [2] estimator with hyperparameter $C=10$ given its robustness with natural image data. Our experiments indicate that ID estimation based on Fisher separability of data works well on both synthetic and real-life datasets, and is more reliable under higher degrees of imbalance.
Using Class-wise ID. Our results suggest that ID-based mitigation works best when used as a resampling or a reweighting strategy. Moreover, although our results show that ID estimation is robust to noise in the data, it can be negatively affected by the number of samples in a class. As a fallback option in such cases, we suggest using the linear-scaling mean ID fallback introduced in Section 6.5.
7.3 Limitations and Future Work
ID-based mitigation shows promise with long-tailed datasets, however, accurately estimating the ID of samples with very low sample sizes remains a challenge. We believe that with better ID estimation tools in the future, our method can improved in this limiting aspect.
Another possible direction in future work regarding ID-based mitigation is to further test the robustness of our proposed method through even broader evaluation across diverse datasets, different modalities and real-world settings. Similarly, exploring extensions that improve the interpretability of ID-based mitigation or reduce the computation cost is a promising idea to increase the practicality of our proposed approach.
@@ Acknowledgments
Appendix A Proof Sketch for the Theoretical Motivation
Let the total support of the data $\mathcal{X}$ be the union of $K$ disjoint compact Riemannian sub-manifolds, where each manifold $\mathcal{M}_{k}$ corresponds to a class indexed by $k$:
$\mathcal{X}=\bigcup_{k=1}^{K}\mathcal{M}_{k}.$ (21)
Each class manifold $\mathcal{M}_{k}$ has its own distinct intrinsic dimension $d_{k}$, volume $V_{k}$, and prior probability $\pi_{k}$.
A.1 Decomposition of Generalization Error
We aim to learn a classifier $f:\mathcal{X}\to\{1,\dots,K\}$. The global risk (error) $R(f)$ can be decomposed into the conditional risks for each class manifold:
$R(f)=\sum_{k=1}^{K}\pi_{k}\cdot R_{k}(f),$ (22)
where $R_{k}(f)=\mathbb{E}_{x\sim\mathcal{M}_{k}}[\mathbb{I}(f(x)\neq k)]$ is the generalization error specifically on class $k$.
To achieve a good global performance (and specifically, good performance on the minority class), we must bound the risk $R_{k}(f)$ for every class $k$.
A.2 Covering Number and Manifold Intrinsic Dimensionality
Narayanan & Niyogi [53] showed that the complexity of learning the decision boundary restricted to the manifold $\mathcal{M}_{k}$ is governed by the covering number $\mathcal{N}(\epsilon,\mathcal{H}|_{\mathcal{M}_{k}})$ of that specific manifold. The paper states that the number of balls of radius $\epsilon$ required to cover the space of smooth boundaries on a $d$-dimensional manifold is related to the manifold’s volume and dimension:
Lemma 1 (Component-wise Complexity).
The metric entropy of the hypothesis class restricted to the $k$-th manifold $\mathcal{M}_{k}$ scales with its specific intrinsic dimension $d_{k}$:
$\log\mathcal{N}(\epsilon,\mathcal{H}|_{\mathcal{M}_{k}})\propto V_{k}\cdot\left(\frac{1}{\epsilon}\right)^{d_{k}}.$ (23)
A.3 Local Metric Entropy and Sample Complexity
Using standard Uniform Convergence bounds (e.g., VC or Rademacher), the number of samples $n_{k}$ drawn from class $k$ required to bound the error $R_{k}(f)\leq\epsilon$ with high probability is:
$n_{k}\gtrsim\Omega\left(\left(\frac{1}{\epsilon}\right)^{d_{k}}\right).$ (24)
Appendix B Extended Literature
B.1 Literature on Long-Tailed Visual Recognition